(Calculating an APR and EAR) You're in need of some money fast, and rather than ask your folks for help, you've decided to look into a payday loan. At a payday loan shop right near your school you see that you can borrow $600 and repay $662 in 11 days. What are the APR and the EAR on this payday loan? a. What is the APR on this payday loan? (Round to two decimal places). b. What is the EAR on this payday loan? (Round to two decimal places).
Answer #1

Answer:

given

Amount Borrowed L=$600

Repayment amount A=$662

Number of Days N=11 Days

Let r be the rate of interest per day

A=L*(1+r)^N

662=600*(1+r)^11

r=0.90%

Let in a year 360days are there

So APR= 360*0.90%=323.27%

EAR =(1+0.90%)^360-1=2398.48%